I tried to do this problem by myself also, but I am getting a negative number. I am setting them up wrong, but not sure how to do it the right way.
Marvelous Marilyn wants to combine a 30% alcohol solution with a 50% alcohol solution to get 500 mg of an 80% alcohol solution. Please help her accomplish this task.
I set it up like this:
.30X + .50(500-X) = .80(500)
I am getting a negative number, so I know I did it wrong.
Thanks

Answer by Paul(988)

0.8(500)-0.3x=0.5(500+x)
400-0.3x=250+0.5x
0.8x=150
x=187.5
500-187.5 = 312.5

Hence, about 187.5Acholol for 30% must be added and about 312.5 acholol for the 50% must be added.
Paul.